OpenStudy (anonymous):

if one zero of the quadratic polynomial p(x)=4x^2-8kx-9 is negative of the other,find the value of k.

OpenStudy (shubhamsrg):

sum of roots = -b/a here sum of roots =0 (given indirectly) b=-8k a=4 just make this substitution and find value of k

OpenStudy (anonymous):

@Diyadiya cn u do d resttt

OpenStudy (shubhamsrg):

lol..ofcorse @Diyadiya can.. why not you?? @Parvathysubhash

OpenStudy (anonymous):

heehee

OpenStudy (shubhamsrg):

o.O comeon,,you cant be dumber than me!! :|

OpenStudy (anonymous):

noooo otherwise u do :P

OpenStudy (shubhamsrg):

where are you stuck,,tell me..

OpenStudy (anonymous):

dt i dnt knw u solve d full prob

OpenStudy (shubhamsrg):

hmmm,,the answer which your are trying to extract from me,,you are not gonna get it :D

OpenStudy (shubhamsrg):

in ax^2 + bx + c =0 ,,what is the sum of roots equal to ?

OpenStudy (anonymous):

no prob do it by step by step easy method

OpenStudy (shubhamsrg):

you're gonna do it,,not me,,i'll just be trying to guide you mam

OpenStudy (anonymous):

i won't tlk to u :'(

OpenStudy (shubhamsrg):

hmm,,nevermind,,not my loss atleast.. :|

OpenStudy (anonymous):

k

OpenStudy (shubhamsrg):

why dont you wanna learn?

OpenStudy (anonymous):

i am busy wid another probs yar i wanna submit dese probs at 3 o clock

OpenStudy (shubhamsrg):

no sympathy from my side,,lol.. :D

OpenStudy (anonymous):

plzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz